# Unravelling the asphericities in the explosion and multi-faceted circumstellar matter of SN 2023ixf

### v1

@@ 8.4.2 Photospheric phase polarization evolution: Peak B
A shift in position angle is observed in the imaging polarization evolution at 4.4 d, suggesting a transition in the primary source of continuum luminosity, which was earlier dominated by interaction with the dense CSM. The aforementioned epoch aligns with the peak ionization and UV flux (refer to Table 1 for the timeline); this PA change indicates an alternate source of polarized light. Moving forward, we observe a polarization rise to another peak at 6.4 d (Peak B; 0.54%). The polarization is likely a result of the recombination emission from an aspherical hydrogen envelope or additional luminosity from interaction with an asymmetric CSM. Following Peak B, we see a steady decline in polarization while the PA continued evolving slowly across the early plateau phase, falling to as low as 0.11% on 22 d. Our next epoch of imaging polarization at 28 d sandwiched by two epochs of spectropolarimetric observations (at 27.3 d and 30.9 d), which showed an increase in polarization with the PA changing 50∘ from the polarization minima at 22 d. The early evolution of SN 2023ixf until the middle of the plateau phase showcases the dominant axis in the CSM-outer ejecta, which is illustrated in the q-u plane in Figure 12. Such a preferred alignment onto a dominant axis is a marker of global axisymmetry (Tanaka et al., 2017). Since Peak A and Peak B arise from a roughly axisymmetric geometry of the CSM and outer ejecta, the SN is viewed from an angle of inclination of 60 degrees to the plane of the sky. Larger viewing angles will require more significant asymmetries since the polarization will scale as , where is the angle measured from the pole (Hoflich, 1991; Wang et al., 2001). We explore the origin of Peak B from two potential scenarios outlined below:
@@
Asphericity in the hydrogen envelope: We see the emergence of the broad P-Cygni in SN 2023ixf at 7 d, which coincides closely with the epoch of Peak B in polarization. This likely indicates that the polarization might arise due to the asphericity in the hydrogen envelope of the SN ejecta since the photosphere would reside in the hydrogen envelope during the photospheric phase. Additionally, we observe the appearance of a broad high-velocity absorption feature blueward of H and H, alongside an absorption feature associated with their P-Cygni profiles. The PA change of 90 degrees from Peak A to Peak B emphasizes that the faster-moving ejecta (Peak B) lies perpendicular to the equatorial CSM (Peak A). The switch in PA was also promptly observed between 3.5 and 4.6 d in the spectropolarimetric observations of (Vasylyev et al., 2023) wherein the PA of H at 3.5 d becomes the PA of the continuum at 4.6 d. This likely means that such a transition occurred earlier for Balmer line emission than for the continuum along the CSM-outer ejecta dominant axis. We hence firmly believe that the source of the dual broad Balmer absorption features and polarization Peak B arises from the aspherical shock front in SN 2023ixf, induced by its interaction with the equatorial dense CSM. We explain this further in the following subsection.
@@ 8.4.3 Aspherical shock structure in SN 2023ixf
An aspherical shock structure is crucial for prolonging the survival time of dust grains, resulting in time-varying extinction as was presented in the pre-discovery photometry of SN 2023ixf, 1 day from the explosion (Li et al., 2023). In addition, the aspherical shock front is consistent with a prolonged and weak initial breakout flash in SN 2023ixf and a rising luminosity produced as ejecta flung out after breakout, expanding, and cooling. The asymmetric nature of the dense CSM also explains the low X-ray luminosity of SN 2023ixf in comparison to interacting SNe II (Panjkov et al., 2023), in tandem with the highly asymmetric SN 2017hcc, which exhibited a bright optical-IR emission but a weak X-ray emission (Chandra et al., 2022). Panjkov et al. (2023) described that the stacked X-ray Swift-XRT flux arises from a two-component bremsstrahlung emission consisting of a heavily absorbed hotter component and a less-obscured cooler component, which fits the description of an aspherical shock front in SN 2023ixf.
@@
The primary source of optical emission is driven by the shock decelerated by the equatorial CSM ( 8,500 ), forming the primary P-Cygni Balmer profile. In contrast, the X-ray observations by NuSTAR and Chandra at epoch 11 – 13 d likely originate from the faster shock ( 13,500 ) arising from the freely expanding ejecta interacting with the low-density extended CSM, also leading to the HV Balmer absorption. The mass-loss rate estimates from X-ray observations are consistent with our estimate for the extended CSM (see Section 7. However, Chandra et al. (2023) estimated that the radial evolution of X-ray emission past 13 d evolved as is consistent with a wind profile of whereas our hydrodynamical modeling estimated a steeper wind profile of for the extended CSM, indicating a disagreement (see Section 7).
@@ 8.5 Geometry of the Inner Ejecta / He-Core: Peak C
We also observe signs of asymmetry in the early nebular phase spectroscopy in the form of redshifted excess in H and [Ca II] 7291, 7324 at around +5,000 in Figure 9. We observe a dual-peaked axisymmetric profile of the [O I] doublet in SN 2023ixf during the early nebular phase in Figure 9. In the classification of [O I] line profiles in CCSNe, as proposed by Taubenberger et al. (2009), the double-peaked structure of the [O I] lines exhibiting axisymmetric blueshifted and redshifted peaks is associated with a toroidal distribution of oxygen. Such excitation of oxygen is expected from an asymmetric distribution of 56Ni (Gerardy et al., 2000) when observed from approximately 60 to 90 degrees with respect to the symmetry axis (Maeda et al., 2008). The line profile of [O I] doublet in the early nebular phase conforms to that picture, echoing the polarization signature detected during the early nebular phase. The effect of asymmetric distribution of the 56Ni has also been proposed for these deviations from asphericity (Chugai, 2006; Dessart & Hillier, 2011). In the case of SN 2004dj, (Chugai et al., 2005) argued that the bipolar ejection of 56Ni led to the asymmetric excitation of hydrogen within a spherical hydrogen envelope due to the shock propagation not being purely radial. In such a scenario, a rapidly but smoothly changing PA might be expected during the transition from the photospheric phase to the nebular phase, during which 56Ni is uncovered. This could explain the PA rotation observed in SN 2023ixf from 68 d up until 140 d. Such an ejection of 56Ni is hypothesized to stem directly from bipolar explosions in CCSNe resulting from mildly rotating RSG cores (Chugai et al., 2005).
@@ 9 Summary
Hydrodynamical light curve modeling: Using STELLA to perform hydrodynamical light curve modeling of SN 2023ixf, we estimated a progenitor mass of 10 with a radius of 470 and an explosion energy of and 0.06 of . The early UV excess was best modeled by a confined dense CSM spanning from the tip of the progenitor to arising from a mass-loss rate of 10-2 and a density structure. The late-plateau UV excess was modeled by an extended CSM spanning to with a mass-loss rate of 10-4 and a density structure.
Asphericity in CSM, ejecta, and the He-core: The temporal evolution of ISP-corrected -band polarization in SN 2023ixf showed three distinct peaks: Peak A [, ] at 1.4 d, Peak B [ at 6.4 d, ] and Peak C [, ] at 79.2 d. The observations indicate the presence of an asymmetric confined dense CSM, an aspherical shock front, and clumpiness in the low-density extended CSM, along with an aspherical inner He-core in SN 2023ixf. The inference on an aspherical shock front also stems from the dual broad absorption features seen in the blue wing of the Balmer features starting 16 d. Overall, SN 2023ixf displayed two dominant axes, one along the CSM-outer ejecta and the other along the inner ejecta/He-core, showcasing the independent origin of asymmetry during the early and late evolution.
Signs of Molecular CO formation: The flattening in the -band light curve and the attenuation of the red-edge of H post 125 d indicates early onset of molecular CO and hence dust formation in SN 2023ixf similar to SN 2017eaw and SN 1987A.
SN 2023ixf, the second closest CCSNe observed in the 21st century, merits deeper exploration to comprehensively analyze its high-cadence observations across the electromagnetic spectrum. Conducting late-phase nebular spectroscopy holds promise in illuminating the intricate, asymmetric structure of the CSM surrounding SN 2023ixf and providing insights into the properties of its synthesized dust. We anticipate that the aspherical shock structure may manifest in the observations at radio wavelengths, revealing evidence of both freely expanding and decelerated shocks. The interaction of the faster shock with the low-density CSM is expected to generate a rapid rise in the radio light curve, while the interaction with the equatorial dense CSM may result in a slower increase over the subsequent months. These distinct radio signatures prompt further observational follow-up to validate the findings outlined in this article.
@@ acknowledgments
We thank the IAO, Hanle, CREST, and Hosakote staff, who made the observations from HCT possible. The facilities at IAO and CREST are operated by the Indian Institute of Astrophysics, Bangalore. The GROWTH India Telescope (GIT) is a 70-cm telescope with a 0.7-degree field of view, set up by the Indian Institute of Astrophysics (IIA) and the Indian Institute of Technology Bombay (IITB) with funding from Indo-US Science and Technology Forum and the Science and Engineering Research Board, Department of Science and Technology, Government of India. It is located at the Indian Astronomical Observatory (IAO, Hanle). We acknowledge funding by the IITB alumni batch of 1994, which partially supports the operation of the telescope. Telescope technical details are available at GROWTH-India Website999https://sites.google.com/view/growthindia/. This research is based partly on data obtained through the KASTOR (Kanata And Seimei Transient Observation Regime) program. The Seimei telescope at the Okayama Observatory is jointly operated by Kyoto University and the National Astronomical Observatory of Japan (NAOJ), with assistance provided by the Optical and Infrared Synergetic Telescopes for Education and Research (OISTER) program funded by the MEXT of Japan. The authors thank the TriCCS developer team (supported by the JSPS KAKENHI grant Nos. JP18H05223, JP20H00174, and JP20H04736 and by NAOJ Joint Development Research).
TJM is supported by the Grants-in-Aid for Scientific Research of the Japan Society for the Promotion of Science (JP24K00682, JP24H01824, JP21H04997, JP24H00002, JP24H00027, JP24K00668) and the Australian Research Council (ARC) through the ARC’s Discovery Projects funding scheme (project DP240101786). KM acknowledges support from the JSPS KAKENHI grant JP20H00174 and 24H01810. DKS acknowledges the support provided by DST-JSPS under grant number DST/INT/JSPS/P 363/2022. GCA thanks the Indian National Science Academy for support under the INSA Senior Scientist Programme.
Numerical computations were in part performed by the PC cluster at the Center for Computational Astrophysics (CfCA), National Astronomical Observatory of Japan. This research made use of RedPipe101010https://github.com/sPaMFouR/RedPipe (Singh, 2021), an assemblage of data reduction and analysis scripts written by AS. This research also made use of the NASA/IPAC Extragalactic Database (NED111111https://ned.ipac.caltech.edu), which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology.

### v2

@@ 8.4.2 Photospheric phase polarization evolution: Peak B
We observe a shift in the imaging polarization evolution at 4.4 d, suggesting a transition in the primary source of continuum luminosity, which was earlier dominated by interaction with the dense CSM. The aforementioned epoch aligns with the peak ionization and UV flux (refer to Table 1 for the timeline); this PA change indicates an alternate source of polarized light. Moving forward, we observe a polarization rise to another point at 6.4 d (Peak B; 0.54%). The polarization is likely a result of the recombination emission from an aspherical hydrogen envelope or additional luminosity from interaction with an asymmetric CSM. Following Peak B, we see a steady decline in polarization while the PA continued evolving slowly across the early plateau phase, falling to as low as 0.11% on 22 d. The epoch of polarization minima coincides with the changing slope during the plateau phase (). Our next epoch of imaging polarization at 28 d sandwiched by two epochs of spectropolarimetric observations (at 27.3 d and 30.9 d), which showed an increase in polarization with the PA changing 50∘ from the polarization minima at 22 d. The early evolution of SN 2023ixf until the middle of the plateau phase showcases the dominant axis in the CSM-outer ejecta, which is illustrated in the q-u plane in Figure 12. Such a preferred alignment onto a dominant axis is a marker of global axisymmetry (Tanaka et al., 2017). Since Point A and Peak B arise from a roughly axisymmetric geometry of the CSM and outer ejecta, the SN is viewed from an angle of inclination of 60 degrees to the plane of the sky. Larger viewing angles will require more significant asymmetries since the polarization will scale as , where is the angle measured from the pole (Hoflich, 1991; Wang et al., 2001). We explore the origin of Peak B from two potential scenarios outlined below:
@@
Asphericity in the hydrogen envelope: We see the emergence of the broad P-Cygni in SN 2023ixf at 7 d, which coincides closely with the epoch of Peak B in polarization. This likely indicates that the polarization might arise due to the asphericity in the hydrogen envelope of the SN ejecta since the photosphere would reside in the hydrogen envelope during the photospheric phase. Additionally, we observe the appearance of a broad high-velocity absorption feature blueward of H and H, alongside an absorption feature associated with their P-Cygni profiles. The PA change of 90 degrees from Point A to Peak B emphasizes that the faster-moving ejecta (Peak B) lies perpendicular to the equatorial CSM (Point A). The switch in PA was also promptly observed between 3.5 and 4.6 d in the spectropolarimetric observations of (Vasylyev et al., 2023) wherein the PA of H at 3.5 d becomes the PA of the continuum at 4.6 d. This likely means that such a transition occurred earlier for Balmer line emission than for the continuum along the CSM-outer ejecta dominant axis. We hence firmly believe that the source of the dual broad Balmer absorption features and polarization Peak B arises from the aspherical shock front in SN 2023ixf, induced by its interaction with the equatorial dense CSM. We explain this further in the following subsection.
@@ 8.4.3 Aspherical shock structure in SN 2023ixf
An aspherical shock structure is crucial for prolonging the survival time of dust grains, resulting in time-varying extinction as was presented in the pre-discovery photometry of SN 2023ixf, 1 day from the explosion (Li et al., 2024). In addition, the aspherical shock front is consistent with a prolonged and weak initial breakout flash in SN 2023ixf and a rising luminosity produced as ejecta flung out after breakout, expanding, and cooling. The asymmetric nature of the dense CSM also explains the low X-ray luminosity of SN 2023ixf in comparison to interacting SNe II (Panjkov et al., 2023), in tandem with the highly asymmetric SN 2017hcc, which exhibited a bright optical-IR emission but a weak X-ray emission (Chandra et al., 2022). Panjkov et al. (2023) described that the stacked X-ray Swift-XRT flux arises from a two-component bremsstrahlung emission consisting of a heavily absorbed hotter component and a less-obscured cooler component, which fits the description of an aspherical shock front in SN 2023ixf.
@@
The primary source of optical emission is driven by the shock decelerated by the equatorial CSM ( 8,500 ), forming the primary P-Cygni Balmer profile. In contrast, the X-ray observations by NuSTAR and Chandra at epoch 11 – 13 d likely originate from the faster shock ( 13,500 ) arising from the freely expanding ejecta interacting with the low-density extended CSM, also leading to the HV Balmer absorption. A weak-synchrotron emission arising from interaction with the low-density extended CSM (Hiramatsu et al., 2023) is consistent with the lack of detections in mm-wavelengths by (Berger et al., 2023) post 4 days. The mass-loss rate estimates from X-ray observations are consistent with our estimate for the extended CSM (see Section 7. However, Chandra et al. (2023) estimated that the radial evolution of X-ray emission past 13 d evolved as is consistent with a wind profile of whereas our hydrodynamical modeling estimated a steeper wind profile of for the extended CSM, indicating a slight disagreement (see Section 7). However, it is important to acknowledge that neither the 1-D hydrodynamical modeling nor a single component X-ray emission fully captures the complexities of asymmetry in the CSM/ejecta of SN 2023ixf as observed through its polarization evolution. Initial attempts towards a two-component X-ray emission from (Panjkov et al., 2023) attribute it to the asphericity in SN 2023ixf, potentially explaining its underluminous nature. Future studies with 2-D/3-D hydrodynamical modeling with the wealth of information on SN 2023ixf will be necessary to address these details and further our understanding of its progenitor system.
@@ 8.5 Geometry of the Inner Ejecta / He-Core: Peak C
We observe signs of asymmetry in the early nebular phase spectroscopy in the form of redshifted excess in H and [Ca II] 7291, 7324 at around +5,000 in Figure 9. Additionally, a dual-peaked axisymmetric profile of the [O I] doublet, blueshifted by 1500 , is also observed in SN 2023ixf during the early nebular phase in Figure 9. It is important to note that these signatures of asymmetry are distinct from the blueshifted excess seen close to the emission peak of the line profiles of H and [Ca II] 7291, 7324 which is likely a result of dust formation (see Figure 9). In the classification of [O I] line profiles in CCSNe, as proposed by Taubenberger et al. (2009), the double-peaked structure of the [O I] lines exhibiting axisymmetric blueshifted and redshifted peaks is associated with a toroidal distribution of oxygen. Such excitation of oxygen is expected from an asymmetric distribution of 56Ni (Gerardy et al., 2000) when observed from approximately 60 to 90 degrees with respect to the symmetry axis (Maeda et al., 2008; Fang et al., 2024). The line profile of [O I] doublet in the early nebular phase conforms to that picture, echoing the polarization signature detected during the early nebular phase. The effect of asymmetric distribution of the 56Ni has also been proposed for these deviations from asphericity (Chugai, 2006; Dessart & Hillier, 2011). In the case of SN 2004dj, (Chugai et al., 2005) argued that the bipolar ejection of 56Ni led to the asymmetric excitation of hydrogen within a spherical hydrogen envelope due to the shock propagation not being purely radial. In such a scenario, a rapidly but smoothly changing PA might be expected during the transition from the photospheric phase to the nebular phase, during which 56Ni is uncovered. This could explain the PA rotation observed in SN 2023ixf from 68 d up until 140 d. Such an ejection of 56Ni is hypothesized to stem directly from bipolar explosions in CCSNe resulting from mildly rotating RSG cores (Chugai et al., 2005).
@@ 9 Summary
Hydrodynamical light curve modeling: Using STELLA to perform hydrodynamical light curve modeling of SN 2023ixf, we estimated a progenitor mass of 10 with a radius of 470 and an explosion energy of and 0.06 of . The inferred properties are not a unique solution since degeneracies exist in the modeling. The early UV excess was best modeled by a confined dense CSM spanning from the tip of the progenitor to arising from a mass-loss rate of 10-2 and a density structure. The late-plateau UV excess was modeled by an extended CSM spanning to with a mass-loss rate of 10-4 and a density structure. A wind-velocity of 10 was assumed for computing the mass-loss rates. The presence of the two-zone CSM in SN 2023ixf with 2 orders of magnitude difference in mass-loss rate suggests an increased mass-loss rate of the progenitor as it neared its explosion.
Asphericity in CSM, ejecta, and the He-core: The temporal evolution of ISP-corrected -band polarization in SN 2023ixf showed three distinct peaks: Point A [, ] at 1.4 d, Peak B [ at 6.4 d, ] and Peak C [, ] at 79.2 d. The observations indicate the presence of an asymmetric confined dense CSM, an aspherical shock front, and clumpiness in the low-density extended CSM, along with an aspherical inner He-core in SN 2023ixf. The inference on an aspherical shock front also stems from the dual broad absorption features seen in the blue wing of the Balmer features starting 16 d. Overall, SN 2023ixf displayed two dominant axes, one along the CSM-outer ejecta and the other along the inner ejecta/He-core, showcasing the independent origin of asymmetry during the early and late evolution.
Signs of Molecular CO / Dust formation: The flattening in the -band light curve and the attenuation of the red-edge of H post 125 d indicates early onset of molecular CO and dust formation in SN 2023ixf similar to SN 2017eaw and SN 1987A. However, our dataset lacks an NIR spectrum during the nebular phase to confirm the presence of molecular CO.
SN 2023ixf, the third closest CCSNe observed in the 21st century, merits deeper exploration to comprehensively analyze its high-cadence observations across the electromagnetic spectrum. Conducting late-phase nebular spectroscopy holds promise in illuminating the intricate, asymmetric structure of the CSM surrounding SN 2023ixf and providing insights into the properties of its synthesized dust. We anticipate that the aspherical shock structure may manifest in the observations at radio wavelengths, revealing evidence of both freely expanding and decelerated shocks. The interaction of the faster shock with the low-density CSM is expected to generate a rapid rise in the radio light curve, while the interaction with the equatorial dense CSM may result in a slower increase over the subsequent months. These distinct radio signatures prompt further observational follow-up to validate the findings outlined in this article.
@@ acknowledgments
We thank the anonymous referee for their comments which helped in the readability of the manuscript. We thank the IAO, Hanle, CREST, and Hosakote staff, who made the observations from HCT possible. The facilities at IAO and CREST are operated by the Indian Institute of Astrophysics, Bangalore. The GROWTH India Telescope (GIT) is a 70-cm telescope with a 0.7-degree field of view, set up by the Indian Institute of Astrophysics (IIA) and the Indian Institute of Technology Bombay (IITB) with funding from Indo-US Science and Technology Forum and the Science and Engineering Research Board, Department of Science and Technology, Government of India. It is located at the Indian Astronomical Observatory (IAO, Hanle). We acknowledge funding by the IITB alumni batch of 1994, which partially supports the operation of the telescope. Telescope technical details are available at GROWTH-India Website101010https://sites.google.com/view/growthindia/. This research is based partly on data obtained through the KASTOR (Kanata And Seimei Transient Observation Regime) program. The Seimei telescope at the Okayama Observatory is jointly operated by Kyoto University and the National Astronomical Observatory of Japan (NAOJ), with assistance provided by the Optical and Infrared Synergetic Telescopes for Education and Research (OISTER) program funded by the MEXT of Japan. The authors thank the TriCCS developer team (supported by the JSPS KAKENHI grant Nos. JP18H05223, JP20H00174, and JP20H04736 and by NAOJ Joint Development Research).
TJM is supported by the Grants-in-Aid for Scientific Research of the Japan Society for the Promotion of Science (JP24K00682, JP24H01824, JP21H04997, JP24H00002, JP24H00027, JP24K00668) and the Australian Research Council (ARC) through the ARC’s Discovery Projects funding scheme (project DP240101786). KM acknowledges support from the JSPS KAKENHI grant JP20H00174 and JP24H01810. DKS acknowledges the support provided by DST-JSPS under grant number DST/INT/JSPS/P 363/2022. GCA thanks the Indian National Science Academy for support under the INSA Senior Scientist Programme. BK acknowledges support from the “Science & Technology Champion Project” (202005AB160002) and from two “Team Projects” – the “Top Team” (202305AT350002) and the “Innovation Team” (202105AE160021), all funded by the “Yunnan Revitalization Talent Support Program”.
Numerical computations were in part performed by the PC cluster at the Center for Computational Astrophysics (CfCA), National Astronomical Observatory of Japan. This research made use of RedPipe111111https://github.com/sPaMFouR/RedPipe (Singh, 2021), an assemblage of data reduction and analysis scripts written by AS. This research also made use of the NASA/IPAC Extragalactic Database (NED121212https://ned.ipac.caltech.edu), which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology.